OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (3a^2)(2a^-1)^2 ________________ (15b^3)(2b^-2)

OpenStudy (anonymous):

\[{{(3a²)(2a^{-1})^2 \over (15b³)(2b^{-2})}}={{3a^2 \times 4a^{-2}} \over 30b}={12a^0 \over 30b} = {2 \over 5b}\]